A Bibliography of Publications by, and about, Augusta Ada King, Countess of Lovelace

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 May 2024 Version 1.23

Abstract

This bibliography records publications of Augusta Ada King, Countess of Lovelace.

Title word cross-reference

$149.50 [Gra94a, Gra94b]. $21.50 [Mad86]. $25 [Yon75a, Yon75b]. $25.00 [Hym78a, Hym78b]. $3.95 [LY79a, LY79b]. $9.95 [Par89a, Par89b]. ème [Kno88a, Kno88b].

-1- [Gor83].

1 [Lov53]. 1847 [Kno88a, Kno88b]. 19th [Gra94a, Gra94b].

2 [BL53]. 2000 [Min00]. 2001 [Coc02]. 2015 [Flo16]. 21st [Too98].

6. [Min00]. 6./7.Juli [Min00].

7.Juli [Min00]. 75pp [LY79a, LY79b]. 75th [JL11].

95 [Whe95, Whe97].

A. [Gra94a, Gra94b]. Académie [Mah78a, Mah78b]. Account [Fra81]. activities [Per93]. Ada [ASS87, HG79, Mad86, Mar88, Min00, Rap87, Sol94, Ste89a, Ste88, WK05, Ang05, Ano11, Ano15, Ano21, Aus07, Bau86, CA14a, CA15, Chi17, Cox19, Elw75, Ess14a, Ess14b, EB19, Fis20, Flo16, Fre96, FFT^{+}03, FF03b, Gre21, HP15, HR16, Hey17, HMR17a, Hoo12, Jam81, JM83, KT99, KCB11, KM12, Lea20, Lee00, LK01, Lov12, Mar18, Men21a, Men21b, Moo77b, Moo77a, Mor13, QPEUR^{+}12, Rin15, Sch04, Sey18, Spi03, Ste85, SBK04, Swa11, Swa13, Too91, Too92, Too95, Too96, Too98, Wad94, Whe95, Whe97, Win98, WK05, Har78, Hym78a, Hym78b, Mar79, Mar88, Par89a, Par89b, Tee87]. Ada-Lovelace-Projekt [Min00]. Ada-Lovelace-Projekts [QPEUR^{+}12]. Ada-Lovelace-Schriftenreihe [Ano11]. Addition [Bab43]. Adrian [Fis20]. Adventures [Pad15, Sus11]. After [Jam81]. Age [Ess04, Ess14a, Ess14b, HR16, Too98, SBK04]. Agusta [Men21a]. Air [Sus11]. Alan [Sus11, Swa13]. Albert [GW94]. algebra [Gra94a, Gra94b]. Algorithm [Ess14a]. Almanack [Per79]. am [Min00, QPEUR^{+}12, SBK04]. Amy [GW94]. analitica [For75]. analyst [Too91, Too96]. Analytical [CK89, For75, Fre96, Lov43b, Lov15, Bab43, Lov43a, Lov89, Men21a, Men21b, Pad17]. Analytique [Men43]. ancient [LY79a, LY79b]. Anecdotes [SHH81]. Anfang [SBK04]. angle [Mah78a, Mah78b]. Annabella [Elw75, Sey18]. Année [Kno88a, Kno88b]. Anniversary [Smi71]. Annotated [Ogi86]. Announced [Coc02]. Antiquity [Ogi86, Ali86]. Appendix [BL53, Lov53]. appraisal [HMR17b]. Archon [Mad86]. Argenteuil [GW94]. aristocracy [Lew95]. Arlettte [Mah78a, Mah78b]. Art [Sus11]. Artificial [Moo03, Aus07]. assemble [HP15]. Augusta [Lee00, SBK04, Elw75, Rap87]. Augustin [GW94]. Augustin-Louis [GW94]. Ausbildung [QPEUR^{+}12]. Avril [Kno88a, Kno88b]. Award [Coc02].

Babbage [For75, SHH81, CK89, Fra81, FF03a, FF15, HH80, HG79, KT99, Lov43b, Men43, Men21b, Pad15, Pad17, Sus11, Swa01a, Swa13, Swa17, Wil77, For75]. Back [DeF22]. balance [Tho19]. Band [Eva18]. Barbiche [Mah78a, Mah78b]. Basel [Gra94a, Gra94b]. Baum [Mad86, Mar88, Ste89a, Ste88]. Beaune [Mah78a, Mah78b]. beginning [SBK04]. Behind [Lov02]. beim [QPEUR^{+}12]. Beispiel [QPEUR^{+}12]. Bernard [Mah78a, Mah78b]. berufliche [QPEUR^{+}12]. Better [BBF03, Tho19]. between [KT99]. Bibliography [Ogi86]. Bicentenary [Flo16]. Biographical [Ogi86]. Biographies [Cor87, Wei88]. biography [Pad17]. Birkhäuser [Gra94a, Gra94b]. Birth [Ess04]. Birthday [JL11]. Blanchard [GW94]. Block [Coc02]. BLOG [HMB22]. bodily [Win98]. Book [Fis20, For75, GW94, Gra94a, Gra94b, Har78, Hym78a, Hym78b, Kno88a, Kno88b, LY79a, LY79b, Lea20, Lov02, Mad86, Mah78a, Mah78b, Mar79, Mar88, Par89a, Par89b, Spi03, Ste89a, Ste88, Sus11, Tee87, Yon75a, Yon75b]. Books [Mad86]. Books/The [Mad86]. Boston [Gra94a, Gra94b]. Brain [Lov02, Hoo12]. Brian [JL11]. Bride [Woo99]. Brief [Bow53a]. Britain [Bra89, Sum14]. British [Gra78]. Broad [Eva18]. Bruce [Sus11]. BSHM [Flo16]. BSHM/Gresham [Flo16]. Build [Swa01a]. Byron [Ang05, FFT^{+}03, FF03b, Har78, Hym78a, Hym78b, Mar79, Too96, Wad94, ASS87, Aus07, Bau86, Elw75, Ess14a, Ess14b, Gor83, Mar88, Moo77b, Moo77a, Sey18, Sol94, Too92, Woo99, WK05, Mad86, Ste89a, Ste88]. Byronic [Mad86]. Byrons [WK05].

CACM [HMB22]. Calculating [ASS87, Bau86, Mad86, Mar88, Ste89a, Ste88]. Calculations [DeF22]. calculus [Win98]. Cambridge [Par89a, Par89b, Yon75a, Yon75b]. can [Ano11]. cartoonist [Pad17]. Cauchy [GW94]. Celebrating [Ano15]. Celebration [Flo16]. Century [Ogi86, Too98, Yon75a, Yon75b, Ali86, Gra94a, Gra94b, Lew95]. Chancrin [Mah78a, Mah78b]. chang [Yon75a, Yon75b]. Charles [HG79, Men43, CK89, EB19, HH80, KT99, Lea20, Men21b, SHH81, Swa01a, Swa13]. Ch’in [Yon75a, Yon75b]. China [LY79a, LY79b]. Chinese [LY79a, LY79b, Yon75a, Yon75b]. chiu [Yon75a, Yon75b]. chiu-chang [Yon75a, Yon75b]. Chiu-Shao [Yon75a, Yon75b]. Choix [GW94]. Christopher [Fis20]. Classic [Lew21]. Coding [Tho19]. collaboration [KT99]. Collections [Mah78a, Mah78b]. College [Flo16, Flo16]. Collegiate [Coc02]. Complexities [CL05]. compliance [Sum14]. Components [Lev15]. Computable [Zen13]. Computadoras [Sol94]. Computation [Bow53a, Zen13]. Computations [DeF22]. Computer [Cox19, DeF22, Ess14b, EB19, Fis20, Fra81, Gol72, KT99, Lea20, Lew21, LR67, Mar18, Smi71, Swa01a, Tho19, Too92, Too98, Ano15, Gre21, LK01, Sum14, Wad94, Wil77]. Computers [ASS87, Gon24, Sol94]. Computing [Bow53b, CA15, HR16, JL11, Par89a, Par89b, Swa13, CA14a, Gür02, HP15]. Conditionals [Fra81]. Conn [Mad86]. constraints [Win98]. Context [Ste84, Ste89b]. correspondence [HMR17b]. Costabel [Mah78a, Mah78b]. Countess [Har78, Hym78a, Hym78b, HG79, KT99, Lee00, Mar79, Men21a, Moo77b, Moo77a, JM83]. Created [Isa14, Lew21]. Creation [FF03a, FF15]. Creativity [BBF03, Bod95]. critical [HMR17b]. Cultural [Ste84, Ste89b]. Culture [Bra89]. Cultures [HR16]. Curie [Kno88a, Kno88b].

D. [HG79]. Dahan [GW94]. Dalmedico [GW94]. Data [Cor87, HMB22]. Daughter [Ess14a, Ess14b, Sey18, Too92, Woo99, Hym78a, Hym78b, Lew95, Moo77b, Moo77a, WK05, Har78, Mar79]. Dedicated [JL11]. Defiance [Sum14]. Dependable [JL11]. described [KT99]. Description [Too92]. Design [Gor83]. Despite [Mad86]. DEUGSSM [Kno88a, Kno88b]. d’Histoire [Kno88a, Kno88b, Mah78a, Mah78b]. Diary [Per79]. Dictionary [Cor87, Ogi86]. Difference [Sus11, Fra81, Swa01a]. Digital [Bow53b, Ess14a, Gon24, HR16, Isa14, LR67, Pla98, Swa13]. Discoveries [Lov02]. Discovery [CA14b, Per93]. Disputes [Gon24]. Doctrine [Mah78a, Mah78b]. Does [LR67, Tho19]. Donovan [Sus11]. Doris [Har78, Hym78a, Hym78b, Mar79]. Dorothy [Mar88, Par89a, Par89b, Tee87]. dream [FFT^{+}03, FF03b, Spi03]. Dubbey [HG79]. Dugac [Kno88a, Kno88b]. durchgeführt [Min00]. during [QPEUR^{+}12].

early [HMR17a, Win98]. Eckert [Smi71]. école [GW94]. edge [Hol18]. Edited [Mah78a, Mah78b]. Editions [GW94]. Editor [GW94]. Editors [Gra94a, Gra94b]. education [HMR17a]. Einstein [Hol18]. Electronic [LR67, Smi71]. Elusive [Moo03]. Enchantress [Chi17, Swa01b, Too92, Too98]. Energy [Bra89]. Engine [CK89, For75, Fre96, Bab43, Fra81, Lov43b, Lov43a, Lov89, Lov15, Men21a, Men21b, Pad17, Sus11, Swa01a]. England [LK01, Win98]. English [Lew95]. Entropy [Bra89]. Essays [Bra89, JL11]. Etas [For75]. Evaluation [Min00]. Everything [Mad86]. examination [LY79a, LY79b]. example [QPEUR^{+}12]. excursions [Hol18]. Exploring [Zen13]. Extracts [BL53, Lov53]. Extraordinary [Sus11].

Fachkonferenz [Min00]. family [Elw75, Lee00]. Faster [Bow53b]. Federico [Lov43b]. Female [Ess14b, WW80]. Février [Kno88a, Kno88b]. First [Cox19, DeF22, KT99, Mor13, Swa01a, Too92, Gre21, HMB22, Sol94]. Florimond [Mah78a, Mah78b]. francaise [GW94]. Frank [LY79a, LY79b]. Frau [SBK04]. Frauen [Min00]. Frauenförderung [Min00]. Ftr [GW94]. Future [Lew21].

G [For75]. Geeks [Isa14]. gender [Lew95, Tho19]. Genius [Ess14b]. Geniuses [Isa14]. Gentlemen [Sus11]. German [Ano11, QPEUR^{+}12, SBK04, WK05]. Get [Coc02]. Gibson [Sus11]. Gödel [Hol18]. Gresham [Flo16, Flo16]. Group [Isa14]. Guide [CBSW17].

Hackers [Isa14]. Hamden [Mad86]. Hand [Ess04]. Hand-Loom [Ess04]. Harper [Hym78a, Hym78b]. Hartree [Gon24]. Her [Jam81, Too92, Lee00, Lew95]. heritage [Ali86]. Historic [JL11]. Historical [Cor87, HP15]. History [Bow53a, Gra78, Lov02, Par89a, Par89b, Sus11, Tho19, Ali86, Gür02]. Hochschulrektorenkonferenz [Min00]. Hollings [Fis20]. Hypatia [Ali86].

ich [Ano11]. Idea [Wol16]. Ideas [Lew21, Wol16]. III [Bab43, Lov53]. Illustrated [Mad86]. Impossible [SHH81]. Information [Ess04]. Innovators [HP15, Isa14]. Intelligence [Moo03, Aus07]. Internationale [Mah78a, Mah78b]. Internet [Eva18]. Invented [Men21b]. Invention [CA14b, Pri10]. Inventors [Smi71]. Isaacson [HP15]. Italian [For75].

J [LY79a, LY79b, Mah78a, Mah78b]. J. [HG79]. Jacquard [Ess04]. Joan [Mad86, Mar88, Ste89a, Ste88].

kann [Ano11]. Kao [LY79a, LY79b]. Kevin [Sus11]. knowledge [Win98]. Koblenz [Min00]. Koblenz-Landau [Min00]. Kolmogorov [Gra94a, Gra94b]. Kompas [For75].

L [HG79]. Ladies [Per79]. Lady [Sch04, SHH81, Too96, Che82, Gon24, HH80, Sol94, Ste84, Ste89b, Wad94]. Landau [Min00]. landmark [KT99]. Langley [Har78, Hym78a, Hym78b, Mar79]. Language [Gor83, Jam81, Aus07]. Launched [Ess14a]. League [Sus11]. Led [Ess04]. Legacy [HR16, Ste85, Mar88, Tee87, Par89a, Par89b]. Legitimate [Har78, Mar79, Hym78a, Hym78b, Moo77b, Moo77a]. Letters [Too92]. Libbrecht [Yon75a, Yon75b]. Librairie [Mah78a, Mah78b]. Life [Mar88, Tee87, Par89a, Par89b, Ste85]. lisant [Kno88a, Kno88b]. Literature [ASS87]. Lives [Sey18, Wol16, Per93]. Logic [Pri10, Gra94a, Gra94b]. London [Flo16, Par89a, Par89b]. Look [DeF22]. Loom [Ess04]. Lord [Sus11, Elw75, Ess14a, Ess14b, Sey18, Too92]. Losano [For75]. Louis [GW94]. Love [Che82]. Lovelace [Ano11, Coc02, Fis20, Flo16, Har78, Mar79, Min00, QPEUR^{+}12, SBK04, Sus11, Whe97, WK05, Ang05, Ano11, Ano15, Ano21, BBF03, CA14a, CA15, Che82, Chi17, Cox19, Ess14a, Ess14b, FFT^{+}03, FF03b, FF03a, FF15, Gon24, Gre21, HP15, Hey17, HMR17a, HMR17b, Hoo12, HH80, HG79, JM83, KT99, KCB11, KM12, Lee00, LK01, Lov12, Mar18, Men21a, Men21b, Moo77b, Moo77a, Mor13, Pad15, Pad17, Rap87, Rin15, Sch04, Sey18, Spi03, Ste84, Ste89b, SBK04, SHH81, Swa11, Swa13, Swa17, Too95, Too96, Wad94, Whe95, Win98, WK05, Hym78a, Hym78b, BL53]. Lucasta [Lov96]. Luigi [Lov43b].

M [HG79]. M. [Bab43, Men43]. MA [Par89a, Par89b]. MA/London [Par89a, Par89b]. macchina [For75]. Machine [Men43]. Machines [Bow53b, Pri10]. Made [Eva18]. Makers [Wol16]. Making [Fis20, Mar18]. manuscript [Mah78a, Mah78b]. Marie [Kno88a, Kno88b]. Mario [For75]. Mark [Smi71]. Martin [Fis20]. Massachusetts [Yon75a, Yon75b]. Math [Lew95, You08]. Mathematical [Gra94a, Gra94b, HG79, HMR17a, HMR17b, Pad17]. mathematicians [Per93, Tee81, Tee83]. Mathematics [CL05, Flo16, Gra94a, Gra94b, BS05, Coo88, Ose74, WK05, Yon75a, Yon75b, Gra78]. Mathematik [WK05]. Mathématisations. [GW94]. Mauchly [Smi71]. Meaning [Gon24]. Mechanical [CK89]. mechanization [BS05]. meeting [Flo16]. memoir [Bab43]. mémoire [Kno88a, Kno88b]. memoirs [Bab43, Lov53]. Menabrea [Bab43]. Menebrae [Lov43b]. Mentoring [QPEUR^{+}12]. merely [Hoo12]. metaphysician [Too91, Too96]. Michael [Sus11]. Milan [For75]. Milbanke [Sey18]. Minds [Lov02]. Modelling [Bod95]. modern [SBK04]. Moderne [SBK04]. Moorcock [Sus11]. Moore [Hym78a, Hym78b, HG79, Sus11, Har78, Mar79]. Morgan [HMR17b]. mortal [Hoo12]. Most [EB19, Lea20]. Mr. [Fra81]. much [Tho19]. Museum [Sus11]. My [Hoo12].

N [Gra94a, Gra94b]. name [Lee00]. Named [Jam81]. Nature [Zen13]. naturwissenschaftlichen [Min00]. Neumann [Gol72]. Neuromancer [Sus11]. News [Coc02]. Nineteenth [Ogi86, Ali86, Lew95]. No [Mah78a, Mah78b]. Norrington [Sus11]. Notable [Wol16]. Notation [CK89, Fra81]. ’Notes [FF15, KM12, Lov43b, Lov15, Men21b, Ste84, Men21a, Ste89b, FF03a]. Notions [Men43]. novel [Chi17]. Number [Swa01b, Gra94a, Gra94b]. Numbers [Too92, Too98, Chi17, Per93].

Objection [Gon24]. Occasion [JL11]. October [Flo16]. Oh [Wei88]. once [Tho19]. O’Neill [Sus11]. Ones [Pla98]. Operations [Pri10]. Option [Kno88a, Kno88b]. Origin [Fra81]. original [Mah78a, Mah78b]. Origins [Swa13]. Other [ASS87]. Oxford [Rin15, Sus11].

P [Gra94a, Gra94b, Bab43, Yon75a, Yon75b, For75]. Padua [Sus11]. paper [Lov43b]. Papers [Lew21, BL53]. para [Sol94]. Parallelograms [Lew95]. Paris [GW94, Kno88a, Kno88b, Mah78a, Mah78b]. Part [Bab43]. Partnership [EB19, Lea20]. Pascal [Gol72]. Passages [Gro23]. Passion [ASS87, Bau86, CA14b, Mad86, Mar88, Ste89a, Ste88]. Passionate [EB19, Lea20]. Pathway [Too98]. Pebbles [ASS87]. Pennsylvania [LY79a, LY79b]. People [Wol16]. Personal [Wol16]. Perspectives [Wol16]. Philomaths [WW80]. Philosophique [Mah78a, Mah78b]. Picturing [Pad17]. Pierre [Kno88a, Kno88b, Mah78a, Mah78b]. Pioneer [Hey17, Swa11, KT99, Wil77]. pioneering [Tee81, Tee83]. Pioneers [Lov02, Wei88]. plus [Per93]. Poems [Lov96]. poetess [WK05]. Poetin [WK05]. portrait [FF03b, Spi03]. postwar [Sum14]. pp [Mad86, Par89a, Par89b, Kno88a, Kno88b, Mah78a, Mah78b]. Press [LY79a, LY79b, Mad86, Par89a, Par89b, Yon75a, Yon75b]. Primer [Sol94]. Princess [Lew95]. probability [Gra94a, Gra94b]. Problem [Fra81]. Processing [Cor87]. Profile [Swa11]. Program [Gor83, SHH81, KT99, Sol94]. Programa [Sol94]. Programmer [Cox19, Gre21, HMB22]. Programming [Fra81, Hey17, Pri10, Tho19]. Project [QPEUR^{+}12]. Projekt [Min00]. Projekts [QPEUR^{+}12]. Prophet [Too98]. publication [KT99]. Pythagoras [LY79a, LY79b].

Qualitätskriterium [Min00]. Quest [Swa01a].

Randell [JL11]. re [HMR17b]. re-appraisal [HMR17b]. Real [Coc02, JM83]. Reason [Woo99]. Recursiveness [Fra81]. reflections [HP15]. reply [Bod95]. resulted [KT99]. Reusable [Lev15]. Review [Fis20, For75, GW94, Gra94a, Gra94b, Har78, Hym78a, Hym78b, Kno88a, Kno88b, LY79a, LY79b, Lea20, Lov02, Mad86, Mah78a, Mah78b, Mar79, Mar88, Par89a, Par89b, Spi03, Ste89a, Ste88, Sus11, Tee87, Yon75a, Yon75b]. reviewers [Bod95]. Reviews [ASS87, HG79]. Revolution [Isa14]. Rice [Fis20]. right [LY79a, LY79b]. Romance [Woo99]. Row [Hym78a, Hym78b].

Schriftenreihe [Ano11]. Science [Bra89, CA14b, Lew21, Ogi86, Pri10, Woo99, Ali86, HMB22, You08, Sus11]. Sciences [Kno88a, Kno88b, Mah78a, Mah78b]. Scientific [Lov53, Bab43]. Scientist [Mar18, Fis20]. Secret [Tho19]. Seidel [Kno88a, Kno88b]. Selection [Too92]. seminar [Coo88]. Series [Ano11, Par89a, Par89b]. Services [Coc02]. Shadowy [BS05]. Shao [Yon75a, Yon75b]. Shoe [Mad86]. Shu [Yon75a, Yon75b]. Shu-shu [Yon75a, Yon75b]. Sketch [Lov43a, Lov89, Lov15, Men21a, Men21b]. Society [Gra78]. Software [Lev15]. solide [Mah78a, Mah78b]. Some [Wol16]. Spanish [Sol94]. spanners [BS05]. Standard [Moo03]. Stanford [Coc02]. Started [Ess14b, LR67]. State [LY79a, LY79b]. Steampunk [Sus11, Sus11, Sus11]. Stein [Mar88, Par89a, Par89b, Tee87]. Stephen [Sus11]. Sterling [Sus11]. Stories [CA14b]. Story [Eva18]. String [Mad86]. Student [Coc02]. Studienbarrieren [Min00]. Studiengängen [Min00]. Studienmotivation [Min00]. Studies [Bra89]. suffering [Win98]. Supercomputer [Coc02]. superheroines [HP15]. Swetz [LY79a, LY79b]. Sydney [Sus11]. Symposium [Bow53b, Rin15].

T [LY79a, LY79b]. Taylor [Lov53]. Tech [Mor13]. Technical [Ste84, Ste89b]. Technik [Min00]. technisch [Min00]. technisch-naturwissenschaftlichen [Min00]. Technoculture [Pla98]. Test [BBF03, BBF03, Moo03]. Text [Ste84, Ste89b]. Their [Lov02]. theory [Gra94a, Gra94b, LY79a, LY79b, Gra94a, Gra94b]. Thirteenth [Yon75a, Yon75b]. Thought [Bow53b, KM12, Hol18]. Thread [ASS87]. Thrilling [Pad15, Sus11]. Tochter [WK05]. today [Tho19]. tomorrow [FFT^{+}03, FF03b, Spi03]. Tournament [Coc02]. training [QPEUR^{+}12]. transcribed [Mah78a, Mah78b]. transition [QPEUR^{+}12]. Translation [Men21b, Lov43b, Lov15]. translator [Men21a]. travaux [Mah78a, Mah78b]. triangle [LY79a, LY79b]. Turbulent [Sey18]. Turing [BBF03, CBSW17, Gon24, Moo03, Swa13, Swa17]. tutorial [Whe95, Whe97]. typed [Mah78a, Mah78b].

Übergang [QPEUR^{+}12]. Ulrich

[Yon75a, Yon75b]. Understanding [Zen13]. Universe [Zen13]. Universität [Min00]. Université [Kno88a, Kno88b]. University [LY79a, LY79b]. Untold [Eva18]. Ursula [Fis20].

VI [Kno88a, Kno88b]. Victorian

[CA15, LK01, Bra89, CA14a, HR16, Win98]. videorecording [FFT^{+}03]. Views [Coc02]. vision [BS05]. Visionary [CA15, Mor13, Ano15, CA14a]. Visions [Sum14]. vocational [QPEUR^{+}12]. Vol [CK89, Lov53, Par89a, Par89b, Sus11, Bab43]. Vortrag [Min00]. Vrin [Mah78a, Mah78b].

Wake [Sey18]. walked [Hol18]. Walter [HP15]. want [Ano11]. War [Sus11]. was [Lee00, Ano11, LY79a, LY79b]. Web [Ess04, Coc02]. went [Tho19]. Who [Eva18]. Wife [Sey18]. will [Ano11]. William [Sus11]. Winds [Coc02]. wizard [LK01]. woman [SBK04, Per79]. Women [CL05, Eva18, Flo16, Gür02, Ogi86, Ose74, Per93, Pla98, Tho19, Ali86, Coo88, Tee81, Tee83, Win98, You08]. Work [HG79]. Works [CK89, LR67]. World [Cox19, KT99]. wrong [Tho19].

XII [Bab43]. xix [Par89a, Par89b].

years [Ano15]. York [Hym78a, Hym78b]. Yushkevich [Gra94a, Gra94b].

Z [You08]. Zeros [Pla98].

References

Alic:1986:HHH

[Ali86] Margaret Alic. Hypatia’s heritage: a history of women in science from antiquity through the nineteenth century, volume 720 of Beacon paperback. Beacon Press, Boston, MA, USA, 1986. ISBN 0-8070-6731-8 (paperback), 0-8070-6730-X (hardcover). ix + 230 + 8 pp. LCCN Q130 .A48 1986b. URL http:

//catdir.loc.gov/catdir/
description/hm031/86047510.
html.

Angluin:2005:ABL

[Ang05]  Dana Angluin.  Ada By-

ron Lovelace. In Case and Leggett [CL05], pages 60–

67. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (ebook). LCCN QA27.5 .C66

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2004048843-b.html;
 http://www.loc.gov/catdir/
enhancements/fy0668/2004048843-
d.html;   http://www.loc.
gov/catdir/toc/fy0612/2004048843.
html.

Anonymous:1998:WIW

[Ano11] Anonymous. Was ich will, das kann ich!: Ada-Lovelace-Schriftenreihe. (German) [What I want, I can!: Ada Lovelace Series]. Frauen in Rheinland-Pfalz; Naturwissenschaft und Technik für Frauen. Rheinland-Pfälzisches Mentoring-Netzwerk für Frauen in MINT, Mainz, Germany, 1998–2011. 300 pp. LCCN ????

Anonymous:2015:ALC

[Ano15] Anonymous. Ada Lovelace: Celebrating 200 years of a computer visionary. Oxford University Symposium Web site., 2015. URL http:/

/blogs.bodleian.ox.ac.uk/
adalovelace/symposium/. 10

December 2015 marks the 200th anniversary of the birth of Ada Lovelace, chiefly known for her work on Charles Babbage’s early mechanical general-purpose computer, the Analytical Engine.

Anonymous:2021:AL

[Ano21] Anonymous. Ada Lovelace.

Computer History Museum, Mountain View, CA, USA,

2021.  URL https://www.
computerhistory.org/babbage/
adalovelace/.

Aspray:1987:RPC

[ASS87] William Aspray, K. W. Smillie, and Nancy Stern. Reviews: Pebbles to Computers. The Thread; The Calculating Passion of Ada Byron; other literature. Annals of the History of Computing, 9(2):214– 217, April/June 1987. CODEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1987/pdf/
a2214.pdf;   http://www.
computer.org/annals/an1987/
a2214abs.htm.

Austin:2007:ABL

[Aus07] Andrea Austin. Ada Byron and the language of artificial intelligence. Cambridge Scholars Pub., Newcastle, UK, 2007. 202–215 pp. LCCN ????

Babbage:1843:AMM

[Bab43] Charles Babbage. Addition to the memoir of M. Menabrea on the Analytical Engine. Scientific memoirs, vol. III. Part XII. p. 666. Philosophical Magazine, 23(151):235–239, ???? 1843. CODEN PHMAA4. ISSN 0031-8086. URL http://www.

tandfonline.com/doi/abs/
10.1080/14786444308644724.

Baum:1986:CPA

[Bau86] Joan Baum. The Calculating Passion of Ada Byron. Archon Books, Hamden, CT, USA, 1986. ISBN 0-208-02119-1. xix + 133 + 12 pp. LCCN QA29.L72 B381 1986.

Bringsjord:2003:CTT

[BBF03] Selmer Bringsjord, Paul Bello, and David Ferrucci. Creativity, the Turing Test, and the (better) Lovelace test. In Moor [Moo03], pages 215– 239. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. LCCN Q341 .T87 2003. URL http://

link.springer.com/chapter/
10.1007/978-94-010-0105-2_
12/.

Babbage:1953:AEL

[BL53] Charles Babbage and Ada Augusta Lovelace. Appendix 2: Extracts from the Lovelace Papers. In Bowden [Bow53b], page 409. LCCN QA76.5 .B66.

URL https://archive.org/
details/FasterThanThought.

Quotes from four letters between Charles Babbage and the Countess of Lovelace.

Boden:1995:MCR

[Bod95] Margaret A. Boden. Modelling creativity: reply to reviewers. Artificial Intelligence, 79(1):161–182, November 1995. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0004370295000747.

Bowden:1953:BHC

[Bow53a] Baron Bertram Vivian Bowden.

A brief history of computation. In Faster Than Thought: a Symposium on Digital Computing Machines [Bow53b], pages 3–31. LCCN QA76.5 .B66.

URL https://archive.org/
details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Bowden:1953:FTT

[Bow53b] Baron Bertram Vivian Bowden, editor. Faster Than Thought: a Symposium on Digital Computing Machines. Sir Isaac Pitman and Sons, Ltd., London, UK, 1953. LCCN QA76.5 .B66.

URL https://archive.org/
details/FasterThanThought.

With a foreword by the Right Honourable the Earl of Halsbury.

Brantlinger:1989:EES

[Bra89] Patrick Brantlinger, editor. Energy and Entropy: Science and Culture in Victorian Britain: Essays from Victorian Studies. Indiana University Press, Bloomington, IN, USA, 1989. ISBN 0-253-31928-5. xxii + 352 pp. LCCN Q175.52.G7 E54 1988.

Belanger:2005:SVS

[BS05] Jay Belanger and Dorothy Stein. Shadowy vision: spanners in the mechanization of mathematics. Historia Mathematica, 32(1):76–93, February 2005. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0315086004000217.

Charman-Anderson:2014:ALV

[CA14a] Suw Charman-Anderson. Ada Lovelace: Victorian computing visionary. In A Passion for Science: Stories of Discovery and Invention [CA14b], chapter 8, page ?? ISBN ???? LCCN ???? URL http://

                                        findingada.com/book/.

Charman-Anderson:2014:PSS

[CA14b] Suw Charman-Anderson. A Passion for Science: Stories of Discovery and Invention. ????, ????, 2014. ISBN ???? ???? pp. LCCN ???? URL http://

                                        findingada.com/book/.

Charman-Anderson:2015:ALV

[CA15] Suw Charman-Anderson. Ada Lovelace: Victorian computing visionary. Ada User Journal, 36(1):35–41, March 2015. CODEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Copeland:2017:TG

[CBSW17] B. Jack Copeland, Jonathan Bowen, Mark Sprevak, and Robin Wilson, editors. The Turing Guide. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2017. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). xv + 546 pp. LCCN QA29.T8 C67 2017.

Chesney:1982:LLL

[Che82] Marion Chesney. Love and Lady Lovelace. Hale, London, UK, 1982. ISBN 0-7090-3037-1. 190 pp. LCCN ????

Chiaverini:2017:ENE

[Chi17]  Jennifer Chiaverini.  En-

chantress of numbers: a novel of Ada Lovelace. E. P. Dutton, New York, NY, USA, 2017. ISBN 1-101-98522-4. 436 pp. LCCN

????  URL http://nhcpl.
axis360.baker-taylor.com/
Title?itemid=0020488629.

Campbell-Kelly:1989:WCB-3

[CK89] Martin Campbell-Kelly, editor.

The Works of Charles Babbage, Vol. 3, The Analytical Engine and Mechanical Notation. William Pickering, London, UK, 1989. ISBN 1-85196-503-3, 1-85196-005-8 (set). 253 pp. LCCN ????

Case:2005:CWM

[CL05] Bettye Anne Case and Anne M.

Leggett, editors. Complexities: Women in Mathematics. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-11462-5, 0-691-17109-2, 1-4008-8016-5 (e-book). xix + 412 + 22 pp. LCCN QA27.5 .C66

2005. URL http://www.loc.
gov/catdir/enhancements/
fy0668/2004048843-b.html;

                                         http://www.loc.gov/catdir/
                                        enhancements/fy0668/2004048843-
                                        d.html;   http://www.loc.
                                        gov/catdir/toc/fy0612/2004048843.
                                        html.

Cochran:2002:NVS

[Coc02] Shannon Cochran. News and views: Stanford student winds collegiate tournament; Web services get real; new supercomputer on the block; 2001 Lovelace Award announced. Dr. Dobb’s Journal of Software Tools, 27(7):14, July 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/.

Cooney:1988:SWM

[Coo88] Miriam P. Cooney. A seminar on women and mathematics. Historia Mathematica, 15(4):380–383, November 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0315086088900353.

Cortada:1987:HDD

[Cor87] James W. Cortada. Historical Dictionary of Data Processing: Biographies. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1987. ISBN 0-313-25651-9 (lib. bdg.). xiii + 321 pp. LCCN QA76.15 .C66 1987.

Cox:2019:ALW

[Cox19] Annabelle Cox. Ada Lovelace: the World’s First Computer Programmer. Pen and Sword History, ????, 2019. ISBN 1-5267-3049-9. LCCN ????

DeFranco:2022:CCL

[DeF22] Joanna F. DeFranco. From calculations to computations: A look back at the first computer. Computer, 55(1):82–87, January 2022. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Essinger:2019:CACe

[EB19] James Essinger and Lisa Noel Babbage. Charles and Ada: the Computer’s Most Passionate Partnership. The History Press, Chicago, IL, USA, 2019. ISBN 0-7509-9286-7, 0-7509-9286-7 (e-book). 256 pp. LCCN QA29.L72 E87 2019. URL

http://www.vlebooks.com/
vleweb/product/openreader?
id=none\%26isbn=9780750992862.

Elwin:1975:LBF

[Elw75] Malcolm Elwin. Lord Byron’s family: Annabella, Ada, and Augusta, 1816–1824. John Murray, London, UK, 1975. ISBN 0-7195-3233-7. 252 pp. LCCN ???? Edited from the author’s typescript by Peter Thomson.

Essinger:2004:JWH

[Ess04] James Essinger. Jacquard’s Web: How a Hand-Loom Led to the Birth of the Information Age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2004. ISBN 0-19-280577-0. xi + 302 pp. LCCN QA76.17 .E88

                                        2004. URL http://www.loc.
                                        gov/catdir/enhancements/
                                        fy0620/2004276200-d.html;
                                         http://www.loc.gov/catdir/
                                        enhancements/fy0620/2004276200-
                                        t.html.

Essinger:2014:AAH

[Ess14a] James Essinger. Ada’s Algorithm: How Lord Byron’s Daughter Ada Lovelace Launched the Digital Age. Melville House, Brooklyn, NY, USA, 2014. ISBN 1-61219-408-7 (hardcover), 1-61219-409-5 (e-book). xvi + 254 pp. LCCN QA29.L72 E87 2014.

Essinger:2014:FGH

[Ess14b] James Essinger. A Female Genius: How Ada Lovelace Lord Byron’s Daughter, Started the Computer Age. Gibson Square, London, UK, 2014. ISBN 1-908096-66-7 (hardcover). 256 pp. LCCN QA29.L72 E87 2014.

Evans:2018:BBU

[Eva18]  Claire Lisa Evans.  Broad

Band: the Untold Story of the Women Who Made the Internet. Portfolio, New York, NY, USA, 2018. ISBN 0-7352-1175-2 (hardcover), 0-7352-1176-0 (e-pub). 278 pp. LCCN QA76.2.A2 E93 2018.

Fuegi:2003:LBC

[FF03a] John Fuegi and Jo Francis. Lovelace & Babbage and the creation of the 1843 ‘Notes’. IEEE Annals of the History of Computing, 25(4):16–26, October/December 2003. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (elec-

tronic). URL http://csdl.
computer.org/dl/mags/an/
2003/04/a4016.htm;   http:
//csdl.computer.org/dl/mags/
an/2003/04/a4016.pdf. Women

and gender in the history of computing.

Fuegi:2003:DTP

[FF03b] John Fuegi and Jo Francis. To dream tomorrow: a portrait of Ada Byron Lovelace. Flare Productions, Adelphi, MD, USA, 2003. Video DVD (53m) pp. LCCN ????

Fuegi:2015:LBC

[FF15] J. Fuegi and J. Francis.

Lovelace & Babbage and the creation of the 1843 ’Notes’. Ada User Journal, 36(2):89–98, June 2015. CODEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Fuegi:2003:ABL

[FFT^{+}03] John Fuegi, Jo Francis, Maisie Tomlinson, Steven Crossley, Georgina Terry, et al. Ada Byron Lovelace [videorecording]: to dream tomorrow. Women of power series. Flare Productions, Adelphi, MD, USA, 2003. 1 videodisc (53 min.) pp. LCCN QA29.L72 A333 2003.

Fiss:2020:BRC

[Fis20] Andrew Fiss. Book review:

Christopher Hollings; Ursula Martin; Adrian Rice. Ada Lovelace: The Making of a Computer Scientist. Isis, 111 (1):182–183, March 2020. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

Flood:2016:BGC

[Flo16]  Raymond Flood.  BSHM/

Gresham College meeting: Women in mathematics: A celebration of the bicentenary of Ada Lovelace Gresham College, London, 29 October 2015. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 31(3):245–246, 2016. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2016.1215857.

Forti:1975:BRB

[For75] Umberto Forti. Book review: Babbage, la macchina analitica [(Italian) [Babbage: the Analytical Engine]]: By Mario G. Losano. Milan (Etas Kompas), 1973, 191 p. Historia Mathematica, 2(3):350–353, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086075900853.

Franksen:1981:MBD

[Fra81] Ole Immanuel Franksen. Mr. Babbage, the Difference Engine, and the problem of notation: An account of the origin of recursiveness and conditionals in computer programming.

International Journal of Engineering Science, 19(12):1657– 1694, ???? 1981. CODEN IJE-SAN. ISSN 0020-7225 (print), 1879-2197 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0020722581901580. Special

issue dedicated to Prof. K. Kondo.

Freeman:1996:AAE

[Fre96] Elisabeth Freeman. Ada & the analytical engine. Educom Review, 31(2):??, 1996. CO-DEN EDREEW. ISSN 1045-9146. URL http:

//net.educause.edu/apps/
er/review/reviewArticles/
31240.html.

Goldstine:1972:CPN

[Gol72] Herman Heine Goldstine. The Computer: From Pascal to von Neumann. Princeton University Press, Princeton, NJ, USA, 1972. ISBN 0-691-02367-0 (paperback), 0-691-08104-2. xii + 378 pp. LCCN TK7885.A5 G64 1993.

URL http://www.jstor.org/
stable/j.ctt7rvp0. Second

printing, 1973. Paperback edition 1980. Fifth printing, 1993 with new preface. Reprint 2000 by Books on Demand, Ann Arbor, MI, USA.

Goncalves:2024:LLO

[Gon24] Bernardo Gonçalves. Lady Lovelace’s objection: The Turing–Hartree disputes over the meaning of digital computers, 1946–1951. IEEE Annals of the History of Computing, 46 (1):6–18, January/March 2024. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Gordon:1983:BPD

[Gor83] Michael Gordon. The Byron program design language -1- . ACM SIGADA Ada Letters, 2(4):76–83, January/February 1983. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Gray:1978:BSH

[Gra78] J. J. Gray. The British Society for the History of Mathematics. Historia Mathematica, 5(4):467–468, November 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0315086078902185.

Gray:1994:BRB

[Gra94a] Jeremy Gray. Book review: A. N. Kolmogorov, A. P. Yushkevich, Editors, Mathematics of the 19th century: Mathematical logic, algebra, number theory, probability Theory (1992) Birkhäuser,Boston and Basel 308 $149.50. Historia Mathematica, 21(1):93–94, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Gray:1994:BRK

[Gra94b] Jeremy Gray. Book review: A. N. Kolmogorov, A. P. Yushkevich, Editors, Mathematics of the 19th century: Mathematical logic, algebra, number theory, probability Theory (1992) Birkhäuser,Boston and Basel 308 $149.50. Historia Mathematica, 21(1):93–94, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Gregersen:2021:ALF

[Gre21] E. Gregersen. Ada Lovelace:

The first computer programmer. In Encyclopedia Britannica. Encyclopedia Britannica, 2021. URL https://www.

britannica.com/story/ada-
lovelace-the-first-computer-
programmer.

Groce:2023:Pd

[Gro23] Alex Groce. Passages. ACM SIGSOFT Software Engineering Notes, 48(4):13, October 2023. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617946.3617948.

Gurer:2002:WCH

[Gür02] Denise W. Gürer. Women in computing history. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 34(2):116–120, June 2002. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic). URL

https://www.math.utah.edu/
pub/mirrors/ftp.ira.uka.
de/bibliography/Misc/DBLP/

2002.bib. Women and Computing.

Goldstein:1994:BRA

[GW94] Catherine Goldstein and Paul R. Wolfson. Book review: Amy Dahan Dalmedico, Editor, Mathématisations. Augustin-Louis Cauchy et l’école française (1993) 460 Argenteuil (Editions du Choix) and Paris (Albert Blanchard), 260 Ftr. Historia Mathematica, 21(1):88– 93, February 1994. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic).

Hartman:1978:BRB

[Har78]  Mary S. Hartman.  Book

review: Ada, Countess of Lovelace: Byron’s Legitimate Daughter by Doris Langley Moore. The American Historical Review, 83(4):1011–1012, October 1978. ISSN 0002-8762 (print), 1937-5239 (electronic).

                                        URL http://www.jstor.org/
                                        stable/1867712.

Hey:2017:ALP

[Hey17] Tony Hey. Ada Lovelace programming pioneer. Cambridge University Web site., November 8, 2017. URL http://www.

                                        cambridgeblog.org/2017/11/
                                        ada-lovelace-programming-
                                        pioneer/.

Hyman:1979:RDM

[HG79] Anthony Hyman and N. T. Gridgeman. Reviews: D. L. Moore: Ada, Countess of Lovelace; J. M. Dubbey: The Mathematical Work of Charles Babbage. Annals of the History of Computing, 1(1):75–76, July/September 1979. CODEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1979/pdf/
a1075.pdf;   http://www.
computer.org/annals/an1979/
a1075abs.htm.

Huskey:1980:LLC

[HH80] Velma R. Huskey and Harry D.

Huskey. Lady Lovelace and Charles Babbage. Annals of the History of Computing, 2(4):299–329, October/ December 1980. CODEN AHCOE5. ISSN 0164-1239.

URL http://dlib.computer.
org/an/books/an1980/pdf/
a4299.pdf;   http://www.
computer.org/annals/an1980/
a4299abs.htm.

Hazzan:2022:BDS

[HMB22] Orit Hazzan, Koby Mike, and Herbert Bruderer. BLOG@CACM: Up with data science, and the first programmer. Communications of the ACM, 65(11):10– 11, November 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL https://dl.acm.org/
doi/10.1145/3563107.

Hollings:2017:EME

[HMR17a] Christopher Hollings, Ursula Martin, and Adrian Rice. The early mathematical education of Ada Lovelace. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):221–234, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (elec-

                                        tronic). URL http://www.
                                        tandfonline.com/doi/full/
                                        10.1080/17498430.2017.1325297.

Hollings:2017:LMM

[HMR17b] Christopher Hollings, Ursula Martin, and Adrian Rice. The Lovelace–De Morgan mathematical correspondence: a critical re-appraisal. Historia Mathematica, 44(3):202–231, August 2017. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0315086017300319.

Holt:2018:WEW

[Hol18] Jim Holt. When Einstein walked with Gödel: excursions to the edge of thought. Farrar, Strauss, and Giroux, New York, NY, USA, 2018. ISBN 0-374-14670-5 (hardcover), 0-374-71784-2 (e-book). xi + 368 pp. LCCN PS3608.O4943595 A6 2018.

Hooper:2012:ALM

[Hoo12] Rowan Hooper. Ada Lovelace: My brain is more than merely mortal. New Scientist, 216(2886):29, October 13, 2012. CODEN NWS-CAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0262407912626335.

Haigh:2015:HRI

[HP15] Thomas Haigh and Mark Priestley. Historical reflections: Innovators assemble: Ada Lovelace, Walter Isaacson, and the superheroines of computing. Communications of the ACM, 58(9):20–27, September 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http:

//cacm.acm.org/magazines/
2015/9/191176/fulltext.

Hammerman:2016:AL

[HR16] Robin Hammerman and Andrew L. Russell. Ada’s Legacy: Cultures of Computing from the Victorian to the Digital Age, volume 7 of ACM books. ACM Press, New York, NY 10036, USA, 2016. ISBN 1-970001-51-8 (hardcover), 1-970001-48-8 (paperback), 1-970001-49-6 (ebook), 1-970001-50-X (ePub). ISSN 2374-6777. xii + 249 pp. LCCN QA76.2.L72 H252 2016.

Hyman:1978:BRA

[Hym78a] Anthony Hyman. Book review:

Ada, Countess of Lovelace: Byron’s legitimate daughter: By Doris Langley Moore. New York (Harper & Row). 1977. $25.00. Historia Mathematica, 5(3):366–367, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086078901283.

Hyman:1978:BRB

[Hym78b] Anthony Hyman. Book review: Ada, Countess of Lovelace: Byron’s legitimate daughter: By Doris Langley Moore. New York (Harper & Row). 1977. $25.00. Historia Mathematica, 5(3):366–367, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0315086078901283.

Isaacson:2014:IHG

[Isa14] Walter Isaacson. The Innovators: How a Group of Hackers, Geniuses and Geeks Created the Digital Revolution. Simon and Schuster, New York, NY, USA, 2014. ISBN 1-4711-3879-8 (hardcover), 1-4767-0869-X (cloth), 1-4711-3897-6 (paperback), 1-4104-7497-6 (cloth), 1-4767-0870-3. viii + 542 pp. LCCN QA76.2.A2.

James:1981:ATN

[Jam81] C. J. James. Ada: They named a language after her. Softalk, 2 (??):??–?, November 1981.

Jones:2011:DHC

[JL11] Cliff B. Jones and J. L. (John L.) Lloyd, editors. Dependable and Historic Computing: Essays Dedicated to Brian Randell on the Occasion of His 75th Birthday, volume 6875 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-24540-4, 3-642-24541-2. ISSN 0302-9743. xii + 523 pp. LCCN

QA76.17. URL http://link.
springer.com/10.1007/978-
3-642-24541-1.

James:1983:RAC

[JM83] Carol L. James and Duncan E.

Morrill. The real Ada, countess of Lovelace. ACM SIGSOFT Software Engineering Notes, 8 (1):30–31, January 1983. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Krysa:2011:AL

[KCB11] Joasia Krysa and Carolyn Christov-Bakargiev. Ada Lovelace, volume Documenta (13), [9/6/2012 — 16/9/2012] = 100 Notizen — 100 Gedanken / [Documenta–und–Museum– Fridericianum–Veranstaltungs– GmbH. Artistic dir.: Carolyn Christov–Bakargiev …]; 055 of 100 notes — 100 thoughts. Hatje Cantz, Ostfildern, Germany, 2011. ISBN 3-7757-2904-6, 3-7757-3084-2. 34 pp. LCCN ???? Introd by Joasia Krysa.

Krysa:2012:ALN

[KM12] Joasia Krysa and Luigi Federico Menabrea. Ada Lovelace: 100 Notes, 100 Thought, volume 055. Hatje Cantz Verlag, Kassel, Germany, 2012. ISBN 3-7757-2904-6 (paperback). 34 pp. LCCN QA29.L72 A33 2011.

Knobloch:1988:BRB

[Kno88a] Eberhard Knobloch. Book review: En lisant le mémoire de Seidel de 1847: Par Pierre Dugac. Université Pierre et Marie Curie (Paris VI), DEUGSSM. 2ème Année, Option d’Histoire des Sciences (Février–Avril 1984). Paris. 1984. 29 pp. Historia Mathematica, 15(1):90–97, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086088900596.

Knobloch:1988:BRL

[Kno88b] Eberhard Knobloch. Book review: En lisant le mémoire de Seidel de 1847: Par Pierre Dugac. Université Pierre et Marie Curie (Paris VI), DEUGSSM. 2ème Année, Option d’Histoire des Sciences (Février–Avril 1984). Paris. 1984. 29 pp. Historia Mathematica, 15(1):90–97, February 1988. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086088900596.

Kim:1999:AFC

[KT99] Eugene Eric Kim and Betty Alexandra Toole. Ada and the first computer: The collaboration between Ada, Countess of Lovelace, and computer pioneer Charles Babbage resulted in a landmark publication that described how to program the world’s first computer. Scientific American, 280(5):76–81, May 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Leatherdale:2020:BRC

[Lea20] Dik Leatherdale. Book review: Charles & Ada: The computer’s most passionate partnership. Resurrection: The Journal of the Computer Conservation Society, ?? (89):??, Spring 2020. ISSN 0958-7403. URL https://

computerconservationsociety.
org/resurrection/res89.htm#
e.

Lee:2000:AAC

[Lee00]  J. A. N. Lee.  Ada Au-

gusta, Countess of Lovelace, 1815–1852 what was her family name? IEEE Annals of the History of Computing, 22 (4):72–73, October/December 2000. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel5/
85/19196/00887999.pdf.

Levine:2015:RSC

[Lev15] Trudy Levine. Reusable software components. ACM SIGADA Ada Letters, 35(2): 15–21, August 2015. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Lewis:1995:PPH

[Lew95] Judith S. Lewis. Princess of Parallelograms and her daughter: Math and gender in the nineteenth century English aristocracy. Women’s Studies International Forum, 18 (4):387–394, July/August 1995. CODEN WSINDA. ISSN

                                        ????  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/027753959580030S.

Lewis:2021:ICF

[Lew21] Harry R. Lewis. Ideas That Created the Future: Classic Papers of Computer Science. MIT Press, Cambridge, MA, USA, 2021. ISBN 0-262-04530-3. xxii + 495 pp. LCCN Q124.6-127.2.

Lethbridge:2001:ALC

[LK01] Lucy Lethbridge and Lesley Kerman. Ada Lovelace: the computer wizard of Victorian England. Short, London, UK, 2001. ISBN 1-904095-02-X (paperback). 93 pp. LCCN ????

Lovelace:1843:SAE

[Lov43a] Ada Augusta Lovelace. Sketch of the Analytical Engine. Scientific Memoirs, 3(??):666–731, ???? 1843. Reprinted in [Lov89].

Lovelace:1843:NTL

[Lov43b] Augusta Ada Lovelace. Notes [on translation of Luigi Federico Menebrae’s paper on Babbage’s Analytical Engine]. In Taylor’s Scientific Memoirs, page ????

????, ????, August 1843. LCCN ???? Translation to English, with notes about three times the length of the original French article, of [Men43]. See [Woo99, pages 260–282] and [Swa01a, pages 162–164] for accounts of how those notes were written, and despite vigorous opposition from Charles Babbage [Bab43], finally published in this book. The author was listed only as the initials ‘A. A. L.’, because it was then considered inappropriate for a woman, particularly one of high rank, to author a scientific paper.

Lovelace:1953:AET

[Lov53] Ada Augusta Lovelace. Appendix 1: Extracts from Taylor’s Scientific Memoirs, Vol. III. In Bowden [Bow53b], pages 341–408. LCCN QA76.5 .B66.

URL https://archive.org/
details/FasterThanThought.

Reprint of [Lov43a]. Pages 400– 408 describe the computation of the Bernoulli numbers.

Lovelace:1989:SAE

[Lov89] Augusta Ada Lovelace. Sketch of the Analytical Engine (1843). In Campbell-Kelly [CK89], pages 89–170. ISBN 1-85196-503-3, 1-85196-005-8 (set). LCCN ????

Lovelace:1996:LP

[Lov96] Richard Lovelace. The Lucasta Poems, volume 703 of Project Gutenberg. Project Gutenberg, P.O. Box 2782, Champaign, IL 61825-2782, USA, 1996. URL

                                        ftp://uiarchive.cso.uiuc.
                                        edu/pub/etext/gutenberg/
                                        etext96/lcsta10.zip.

Lovelace:2002:BRB

[Lov02] Robert E. Lovelace. Book review: Minds Behind the Brain: A History of the Pioneers and Their Discoveries. Isis, 93(4): 676–678, December 2002. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

                                        URL http://www.jstor.org/
                                        stable/10.1086/375979.

Lovelace:2012:AL

[Lov12] Ada Lovelace. Ada Lovelace. Hatje Cantz Verlag, Kassel, Germany, 2012. ISBN 3-7757-2904-6, 3-7757-3084-2. 34 pp. LCCN ???? URL http:

                                        //ebooks.ciando.com/book/
                                        index.cfm/bok_id/304705;
                                        http://www.ciando.com/img/
                                        books/width167/3775730842_
                                        k.jpg;   http://www.ciando.
                                        com/pictures/bib/3775730842bib_
                                        t_1.jpg.

Lovelace:2015:NTS

[Lov15] A. A. Lovelace. 1842 notes to the translation of the sketch of The Analytical Engine. Ada User Journal, 36(3):152–181, September 2015. CODEN AU-JOET. ISSN 1381-6551. Bicentennial Ada Lovelace articles.

Lieberman:1967:EDC

[LR67] Henry D. Lieberman and Louis Robinson. The electronic digital computer: How it started, how it works and what it does.

New York Times, ??(??):136– 137, January 9, 1967. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.
proquest.com/docview/118029058?
accountid=14677.

Lay-Yong:1979:BRB

[LY79a] Lam Lay-Yong. Book review:

Was Pythagoras Chinese? An examination of right triangle theory in ancient China: by Frank J. Swetz and T. I. Kao. Pennsylvania (Pennsylvania State University Press). 1977. 75pp. $3.95. Historia Mathematica, 6(2):221–228, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086079900958.

Lay-Yong:1979:BRW

[LY79b] Lam Lay-Yong. Book review:

Was Pythagoras Chinese? An examination of right triangle theory in ancient China: by Frank J. Swetz and T. I. Kao. Pennsylvania (Pennsylvania State University Press). 1977. 75pp. $3.95. Historia Mathematica, 6(2):221–228, May 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086079900958.

Maddox:1986:BRB

[Mad86] Brenda Maddox. Book review:

Byronic, despite everything: The Calculating Passion of Ada Byron, by Joan Baum. Illustrated. 133 pp. Hamden, Conn. Archon Books/The Shoe String Press. $21.50. New York Times, ??(??):519, October 5, 1986. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.
proquest.com/docview/110969134.

Mahoney:1978:BRB

[Mah78a] Micheal S. Mahoney. Book review: Doctrine de l’angle solide: By Florimond de Beaune (1601–1652). Edited by Pierre Costabel, transcribed from the original manuscript by Bernard Barbiche, typed by Arlettte Chancrin. Paris (Librairie Philosophique J. Vrin). (Collections des travaux de l’Académie Internationale d’Histoire des Sciences, No. 19). 1975. 157 pp. Historia Mathematica, 5(3):363–366, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086078901271.

Mahoney:1978:BRD

[Mah78b] Micheal S. Mahoney. Book review: Doctrine de l’angle solide: By Florimond de Beaune (1601–1652). Edited by Pierre Costabel, transcribed from the original manuscript by Bernard Barbiche, typed by Arlettte Chancrin. Paris (Librairie Philosophique J.

Vrin). (Collections des travaux de l’Académie Internationale d’Histoire des Sciences, No. 19). 1975. 157 pp. Historia Mathematica, 5(3):363–366, August 1978. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086078901271.

Marchand:1979:BRB

[Mar79] Leslie A. Marchand. Book review: Ada, Countess of Lovelace: Byron’s Legitimate Daughter, by Doris Langley Moore. Keats–Shelley Journal, 28(5):145–147, ???? 1979. ISSN 0453-4387. URL http://www.

jstor.org/stable/30212847.

Marchand:1988:BRB

[Mar88] Leslie A. Marchand. Book review: Ada: A Life and a Legacy, by Dorothy Stein; The Calculating Passion of Ada Byron, by Joan Baum. Keats– Shelley Journal, 37(3):186–188, ???? 1988. ISSN 0453-4387.

URL http://www.jstor.org/
stable/30209906.

Martin:2018:ALM

[Mar18] Ursula Martin. Ada Lovelace:

the making of a computer scientist. Bodleian Library Oxford University Web site., July 27, 2018. URL http:

//blogs.bodleian.ox.ac.uk/
adalovelace/.

Menabrea:1842:NMA

[Men43] Luigi F. Menabrea. Notions sur la machine analytique de M. Charles Babbage. Bibliothèque Universelle de Genève, 41:352– 376, 1842–1843. CODEN ???? ISSN ???? URL http://

www.fourmilab.ch/babbage/

sketch.html. Translated, with additional notes by Augusta Ada, Countess of Lovelace, as Sketch of the Analytical Engine. Reprinted in [CK89, pages 62– 82]. Goldstine [Gol72, page 19] cites this paper as number 82, reprinted in Taylor’s Scientific Memoirs 111, note A, page 348. Woolley [Woo99, page 400] says it appears in that journal in volume 3, pages 666–731, 1843.

Menabrea:2021:SAEa

[Men21a] L. F. Menabrea. Sketch of the Analytical Engine (1843) / with notes by the translator, Ada Agusta, Countess of Lovelace. In Ideas That Created the Future: Classic Papers of Computer Science [Lew21], chapter 3, pages 9–26. ISBN 0-262-04530-3. LCCN Q124.6-127.2.

Menabrea:2021:SAEb

[Men21b] Luigi Federico Menabrea. Sketch of the Analytical Engine Invented by Charles Babbage: Translation and Notes by Ada Lovelace. Quaternion Books, ????, 2021. 99 pp. LCCN ???? URL

https://quaternionbooks.
com/product/sketch-of-the-
analytical-engine-translation-

and-notes-by-ada-lovelace/
;   https://www.math.utah.
edu/pub/bibnet/authors/b/
babbage-charles.bib.

Minks:2000:SSV

[Min00] Karl-Heinz Minks. Studienmotivation und Studienbarrieren: Vortrag auf der Fachkonferenz “Frauen, Technik, Evaluation — Frauenförderung als Qualitätskriterium in technischnaturwissenschaftlichen Studiengängen”; durchgeführt von der Universität Koblenz-Landau, Ada-Lovelace-Projekt und der Hochschulrektorenkonferenz am 6./7.Juli 2000, volume 2000,8 of Kurzinformation / HIS A. HIS, Hannover, Germany, 2000. 23 pp. LCCN ????

Moore:1977:ACLb

[Moo77a] Doris Langley Moore. Ada, Countess of Lovelace: Byron’s legitimate daughter. Harper & Row, New York, NY, USA, 1977. ISBN 0-06-013012-1. 397 + 6 pp. LCCN CT788.L69 M6 1977; PR4382 .M6 1977.

Moore:1977:ACLa

[Moo77b] Doris Langley-Levy Moore.

Ada, Countess of Lovelace: Byron’s legitimate daughter. J. Murray, London, UK, 1977. ISBN 0-7195-3384-8. [13] + 5– 397 + [12] pp. LCCN PR4382 .M6; DA536.L6 M6. UK£9.50.

Moor:2003:TTE

[Moo03] James H. Moor, editor. The Turing Test: the Elusive Standard of Artificial Intelligence, volume 30 of Studies in cognitive systems. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-1204-7 (hardcover), 1-4020-1205-5 (paperback), 94-010-0105-7 (e-book). ISSN 0924-0780. vi + 273 pp. LCCN Q341 .T87 2003. URL http://www.

springerlink.com/content/
978-94-010-0105-2.

Morais:2013:ALF

[Mor13] Betsy Morais. Ada Lovelace, the first tech visionary. The New Yorker, ??(??):??, October 15, 2013. ISSN 0028-792X. URL http:

//www.newyorker.com/tech/
elements/ada-lovelace-the-
first-tech-visionary.

Ogilvie:1986:WSA

[Ogi86] Marilyn Bailey Ogilvie. Women in Science: Antiquity Through the Nineteenth Century: a Biographical Dictionary with Annotated Bibliography. MIT Press, Cambridge, MA, USA, 1986. ISBN 0-262-15031-X (hardcover), 0-262-65038-X (paperback). xi + 254 pp. LCCN Q141 .O34 1986.

Osen:1974:WM

[Ose74] Lynn M. Osen. Women in mathematics. MIT Press, Cambridge, MA, USA, 1974. ISBN 0-262-15014-X. xii + 185 pp. LCCN QA28 .O83 1974.

Padua:2015:TAL

[Pad15] Sydney Padua. The Thrilling Adventures of Lovelace and Babbage. Pantheon Books, New York, NY, USA, 2015. ISBN 0-307-90827-5 (hardcover), 0-307-90828-3 (e-book). 315 pp. LCCN PN6737.P34 T48 2015.

Padua:2017:PLB

[Pad17]  Sydney Padua.  Picturing

Lovelace, Babbage, and the Analytical Engine: a cartoonist in mathematical biography. BSHM Bulletin: Journal of the British Society for the History of Mathematics, 32(3):214–220, 2017. CODEN ???? ISSN 1749-8430 (print), 1749-8341 (electronic). URL http://www.

tandfonline.com/doi/full/
10.1080/17498430.2017.1315520.

Parshall:1989:BRA

[Par89a] Karen Hunger Parshall. Book review: Ada: a life and a legacy: (MIT Press Series in the History of Computing, Vol. 3). By Dorothy Stein. Cambridge, MA/London (The MIT Press). 1985. xix + 321 pp. $9.95. Historia Mathematica, 16(1):94–95, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086089901055.

Parshall:1989:BRB

[Par89b] Karen Hunger Parshall. Book review: Ada: a life and a legacy: (MIT Press Series in the History of Computing, Vol. 3). By Dorothy Stein. Cambridge, MA/London (The MIT Press). 1985. xix + 321 pp. $9.95. Historia Mathematica, 16(1):94–95, February 1989. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086089901055.

Perl:1979:BDB

[Per79] Teri Perl. The Ladies Diary or Woman’s Almanack, 1704–1841. Historia Mathematica, 6(1):36–53, February 1979. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086079901034.

Perl:1993:WNL

[Per93] Teri Perl. Women and numbers: lives of women mathematicians plus discovery activities. Wide World Publishing/Tetra, San Carlos, CA, USA, 1993. ISBN 0-933174-87-X. 213 pp. LCCN QA28 .P48 1993. US$15.95.

Plant:1998:ZOD

[Pla98] Sadie Plant. Zeros and Ones: Digital Women and the New Technoculture. Fourth Estate, London, UK, 1998. ISBN 0-385-48260-4, 1-85702-386-2 (hardcover), 1-85702-698-5 (paperback). 305 pp. LCCN HQ1233 .P53.

Priestley:2010:SOM

[Pri10] Mark Priestley. A Science of Operations: Machines, Logic and the Invention of Programming. History of computing. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84882-554-4 (hardcover), 1-84882-555-2 (e-book). ix + 341 pp. LCCN QA76.6 .P737 2010.

Quaiser-Pohl:2012:MBU

[QPEUR^{+}12] Claudia Quaiser-Pohl, Martina Endepohls-Ulpe, Ramona Rasic, Tanja Gnosa, and Elisabeth Sander. Mentoring beim Übergang in die berufliche Ausbildung am Beispiel des Ada-Lovelace-Projekts. (German) [Mentoring during the transition to vocational training: the example of Ada Lovelace Project]. Diskurs Kindheitsund Jugendforschung, 7(2): 161–172, 2012. ISSN 1862-5002.

Rappaport:1987:AAL

[Rap87] Karen D. Rappaport. Augusta Ada Lovelace (1815– 1852). In Women of mathematics, pages 135–139. Greenwood, Westport, CT, USA, 1987.

RinoNesin:2015:ALS

[Rin15] Gabriela Asli Rino Nesin. The Ada Lovelace symposium in Oxford. Ada User Journal, 36(4):225–230, December 2015. CODEN AUJOET. ISSN 1381-6551. Bicentennial Ada Lovelace article.

Stein:2004:AAL

[SBK04] Dorothy Stein, Björn Bossmann, and Sabine Kreiner. Ada Augusta Lovelace: eine Frau am Anfang der Moderne. (German) [Ada Augusta Lovelace: a woman at the beginning of the modern age]. Kadmos, Berlin, Germany, second edition, 2004. ISBN 3-931659-64-X. xxiv + 365 pp. LCCN ???? ca. EUR 19.90, ca. EUR AT 20.50, ca. CHF 34.90. URL

http://www.gbv.de/dms/bs/
toc/389886041.pdf.

Schreiber:2004:ALL

[Sch04]  Alfred Schreiber.  Ada,

Lady Lovelace. Mitteilungen der Deutschen Mathematiker-Vereinigung, 12(4):266–267, 2004. ISSN 0947-4471.

Seymour:2018:BWT

[Sey18] Miranda Seymour. In Byron’s Wake: the Turbulent Lives of Lord Byron’s Wife and Daughter: Annabella Milbanke and Ada Lovelace. Simon and Schuster, New York, NY, USA, 2018. ISBN 1-4711-3857-7 (hardcover), 1-4711-3859-3 (ebook). 547 pp. LCCN PR4382 .S49 2018.

Strachey:1981:AIP

[SHH81] C. Strachey, Velma R. Huskey, and Harry D. Huskey. Anecdotes: An impossible program; Charles Babbage and Lady Lovelace. Annals of the History of Computing, 3(4):414–415, October/ December 1981. CODEN AH-COE5. ISSN 0164-1239. URL

http://dlib.computer.org/
an/books/an1981/pdf/a4414.
pdf.

Smith:1971:ECI

[Smi71] William D. Smith. The electronic computer’s inventors: Mauchly and Eckert to mark anniversary. New York Times, ??(??):F7, August 1, 1971. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.
proquest.com/docview/119265082.

SolaecheGalera:1994:LAB

[Sol94] María Cristina Solaeche Galera.

Lady Ada Byron y el primer programa para computadoras. (Spanish) [Lady Ada Byron and the first program for computers]. Divulgaciones Matemáticas, 2(1–2):75– 81, 1994. ISSN 1315-2068.

URL http://www.emis.de/
journals/DM/v2/art5.pdf.

Spicer:2003:TDT

[Spi03] D. Spicer. They dream tomorrow: a portrait of Ada Lovelace [book review]. IEEE Annals of the History of Computing, 25(2):83, April 2003. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Stein:1984:LLN

[Ste84]  Dorothy K. Stein.  Lady

Lovelace’s notes: Technical text and cultural context. Victorian Studies, 28(1):33–67, Autumn 1984. ISSN 0042-5222 (print), 1527-2052 (electronic).

URL http://www.jstor.org/
stable/3826758.

Stein:1985:ALL

[Ste85] Dorothy Stein. Ada, a life and a legacy. MIT Press series in the history of computing. MIT Press, Cambridge, MA, USA, 1985. ISBN 0-262-19242-X (hardcover), 0-262-69116-7 (paperback). xix + 321 + 16 pp. LCCN QA29.L72 S74 1985. US$19.95.

URL https://mitpress.mit.
edu/contributors/dorothy-
stein.

Stern:1988:BRB

[Ste88] Nancy Stern. Book review: The Calculating Passion of Ada Byron, by Joan Baum. American Scientist, 76(1):97, January 1988. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.

jstor.org/stable/27855052.

Stein:1989:BRB

[Ste89a] Dorothy Stein. Book review: The Calculating Passion of Ada Byron, by Joan Baum. Victorian Studies, 32(3):415–416, Spring 1989. ISSN 0042-5222 (print), 1527-2052 (electronic).

URL http://www.jstor.org/
stable/3828502.

Stein:1989:LLN

[Ste89b]  Dorothy K. Stein.  Lady

Lovelace’s notes: Technical text and cultural context. In Brantlinger [Bra89], page ?? ISBN 0-253-31928-5. LCCN Q175.52.G7 E54 1988.

Sumner:2014:DCV

[Sum14] James Sumner. Defiance to compliance: Visions of the computer in postwar Britain. History and Technology, 30(4): 309–333, 2014. CODEN HI-TEE8. ISSN 0734-1512 (print), 1477-2620 (electronic).

Sussman:2011:BRS

[Sus11] Herbert Sussman. Book review: Steampunk at Oxford: Steampunk, by Art Donovan. Steampunk Neuromancer, by William Gibson. The Difference Engine by William Gibson and Bruce Sterling. League of Extraordinary Gentlemen, by Stephen Norrington. The War Lord of the Air by Michael Moorcock. League of Extraordinary Gentlemen, Vol. 1, by Alan Moore and Kevin O’Neill; The Thrilling Adventures of Lovelace and Babbage by Sydney Padua, The Museum of the History of Science. Victorian Literature and Culture, 39(1): 278–284, ???? 2011. ISSN 1060-1503 (print), 1470-1553 (electronic).

Swade:2001:DEC

[Swa01a] Doron Swade. The Difference Engine: Charles Babbage and the Quest to Build the First Computer. Viking, New York, NY, USA, 2001. ISBN 0-670-91020-1, 0-14-200144-9 (paperback). x + 342 + 16 pp. LCCN QA75 .S954 2001. URL http:

//www.gbv.de/dms/bowker/
toc/9780670910205.pdf;
http://www.loc.gov/catdir/
toc/fy034/00068643.html;
http://www.zentralblatt-
math.org/zmath/en/search/?
an=1017.01006.

Swade:2001:EN

[Swa01b] Doron Swade. The enchantress of number. In Swade-DE [Swa01a], chapter 8, pages 155– 171. ISBN 0-670-91020-1, 0-14-200144-9 (paperback). LCCN QA75 .S954 2001. URL http:

//www.gbv.de/dms/bowker/
toc/9780670910205.pdf;
http://www.loc.gov/catdir/
toc/fy034/00068643.html;
http://www.zentralblatt-
math.org/zmath/en/search/?
an=1017.01006.

Swade:2011:PPA

[Swa11] Doron Swade. Pioneer profile: Ada Lovelace. Resurrection: The Computer Conservation Society Journal, ?? (53):31–33, 2011. ISSN 0958-7403. URL https://

computerconservationsociety.
org/resurrection/res53.htm#
g.

Swade:2013:ODC

[Swa13] Doron Swade. Origins of digital computing: Alan Turing, Charles Babbage, and Ada Lovelace. In Zenil [Zen13], pages 23–43. ISBN 981-4374-29-6. LCCN QA267.7 .C676 2013. URL http:/

/www.worldscientific.com/
worldscibooks/10.1142/8306.

Foreword by Roger Penrose.

Swade:2017:TLB

[Swa17] Doron Swade. Turing, Lovelace, and Babbage. In Copeland et al. [CBSW17], chapter 24, pages 249–262. ISBN 0-19-874782-9 (hardcover), 0-19-874783-7 (paperback). LCCN QA29.T8 C67 2017.

Tee:1981:PWM

[Tee81] Garry J. Tee. The pioneering women mathematicians. Mathematical Chronicle (University of Auckland), 10(1–2):31–56, ???? 1981. CODEN MTHCB3. ISSN 0581-1155.

Tee:1983:PWM

[Tee83]  Garry J. Tee.  The pio-

neering women mathematicians. The Mathematical Intelligencer, 5(4):27–36, ???? 1983. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic). URL http://

link.springer.com/article/
10.1007/BF03026507.

Tee:1987:BRB

[Tee87] Garry J. Tee. Book review:

Ada: A Life and a Legacy by Dorothy Stein. Isis, 78 (2):319–320, June 1987. CO-DEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/
stable/231598.

Thompson:2019:SHW

[Tho19] Clive Thompson. The secret history of women in coding: Computer programming once had much better gender balance than it does today. What went wrong? New York Times, February 13, 2019. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https:/

/www.nytimes.com/2019/02/
13/magazine/women-coding-
computer-programming.html.

Toole:1991:AAM

[Too91] Betty Alexandra Toole. Ada, an analyst and a metaphysician. ACM SIGADA Ada Letters, 11(2):60–71, March/April 1991. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Toole:1992:AEN

[Too92] Betty A. (Betty Alexandra) Toole, editor. Ada, the Enchantress of Numbers: a Selection from the Letters of Lord Byron’s Daughter and Her Description of the First Computer. Strawberry Press, Mill Valley, CA, USA, 1992. ISBN 0-912647-09-4. xvi + 439 pp. LCCN QA29.L72 A4 1992. US$29.95. A selection from the letters of Lord Byron’s daughter and her description of the first computer.

Toole:1995:ALI

[Too95] Betty A. Toole. Ada Lovelace. The Mathematics Teacher, 88 (5):435, May 1995. ISSN 0025-5769 (print), 2330-0582 (elec-

tronic). URL http://www.
jstor.org/stable/27969386.

Toole:1996:ABL

[Too96] Betty Alexandra Toole. Ada Byron, Lady Lovelace, an analyst and metaphysician. IEEE Annals of the History of Computing, 18(3):4–12, Fall 1996. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/iel4/
85/11096/00511939.pdf.

Toole:1998:AEN

[Too98] Betty Alexandra Toole. Ada, the Enchantress of Numbers: Prophet of the Computer Age, a Pathway to the 21st Century. Strawberry Press, Mill Valley, CA, USA, revised abridged edition, 1998. ISBN 0-912647-18-3 (paperback). xvii + 323 pp. LCCN QA29.L72 A4 1998.

Wade:1994:ABL

[Wad94] Mary Dodson Wade. Ada Byron Lovelace: the Lady and the computer. People in focus book. Maxwell Macmillan, New York, NY, USA, 1994. ISBN 0-87518-598-3. 128 pp. LCCN QA76.2.L68 W33 1994.

Weiss:1988:BOP

[Wei88]  Eric A. Weiss.  Biogra-

phies: Oh, pioneers! Annals of the History of Computing, 10(4):348–361, October/December 1988. CODEN AHCOE5. ISSN 0164-1239.

                                        URL http://dlib.computer.
                                        org/an/books/an1988/pdf/
                                        a4348.pdf;   http://www.
                                        computer.org/annals/an1988/
                                        a4348abs.htm.

Wheeler:1995:LAT

[Whe95] David A. Wheeler. Lovelace: an Ada 95 tutorial. ACM SIGADA Ada Letters, 15(6): 57–66, November/December 1995. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Wheeler:1997:ALT

[Whe97] David A. Wheeler. Ada 95: the Lovelace tutorial. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 0-387-94801-5 (hardcover). xiv + 292 pp. LCCN QA76.73.A35 W52 1997. URL http://

                                        www.zentralblatt-math.org/
                                        zmath/en/search/?an=0917.
                                        68028.

Wilkes:1977:BCP

[Wil77] M. V. Wilkes. Babbage as a computer pioneer. Historia Mathematica, 4(4):415– 440, November 1977. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0315086077900799.

Winter:1998:CSA

[Win98] Alison Winter. A calculus of suffering: Ada Lovelace and the bodily constraints on women’s knowledge in early Victorian England. In Christopher Lawrence and Steven Shapin, editors, Science incarnate: Historical Embodiments of Natural Knowledge, pages 202–239. University of Chicago Press, Chicago, IL, USA and London, UK, 1998. ISBN 0-226-47012-1. LCCN ????

Woolley:2005:BTA

[WK05] Benjamin Woolley and Katharina Kramer. Byrons Tochter: Ada Lovelace — die Poetin der Mathematik. (German) [Byron’s daughter: Ada Lovelace — the poetess of mathematics], volume 2123 of Aufbau Taschenbücher. Aufbau Taschenbuch Verlag, Berlin, Germany, 2005. ISBN 3-7466-2123-2 (paperback). 453 pp. LCCN ???? EUR 9.95, SFR 18.40. Translation of [Woo99] from English to German by Katharina Kramer.

Wolfram:2016:IMP

[Wol16] Stephen Wolfram. Idea Makers: Personal Perspectives on the Lives and Ideas of Some Notable People. Wolfram Media, Inc., Champaign, IL, USA, 2016. ISBN 1-57955-003-7 (hardcover), 1-57955-005-3 (e-book), 1-57955-011-8. 250 (est.) pp. LCCN Q141 .W678562 2016. URL

http://www.wolfram-media.
com/products/idea-makers.
html.

Woolley:1999:BSR

[Woo99] Benjamin Woolley. The Bride of Science: Romance, Reason, and Byron’s Daughter. Mc-Graw-Hill, New York, NY, USA, 1999. ISBN 0-07-137329-2 (hardcover). viii + 416 pp. LCCN QA29.L72 W66 1999. URL http:/

                                        /www.loc.gov/catdir/bios/
                                        mh041/2001274424.html;
                                        http://www.loc.gov/catdir/
                                        description/mh021/2001274424.
                                        html;   http://www.loc.gov/
                                        catdir/toc/mh021/2001274424.
                                        html.

Wallis:1980:FP

[WW80] Ruth Wallis and Peter Wallis. Female philomaths. Historia Mathematica, 7(1):57– 64, February 1980. CO-DEN HIMADS. ISSN 0315-0860 (print), 1090-249X (elec-

                                        tronic). URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0315086080900646.

Yong:1975:BRB

[Yon75a] Lam Lay Yong. Book review: Chinese mathematics in the Thirteenth Century: the Shu-shu chiu-chang of Ch’in Chiu-Shao: by Ulrich Libbrecht. Cambridge, Massachusetts (MIT Press), 1973. 586 p. $25. Historia Mathematica, 2(3):353–355, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

                                        http://www.sciencedirect.

com/science/article/pii/
0315086075900865.

Yong:1975:BRC

[Yon75b] Lam Lay Yong. Book review: Chinese mathematics in the Thirteenth Century: the Shu-shu chiu-chang of Ch’in Chiu-Shao: by Ulrich Libbrecht. Cambridge, Massachusetts (MIT Press), 1973. 586 p. $25. Historia Mathematica, 2(3):353–355, August 1975. CODEN HIMADS. ISSN 0315-0860 (print), 1090-249X (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0315086075900865.

Yount:2008:ZWS

[You08] Lisa Yount. A to Z of women in science and math. A to Z of women; Facts on File library of world history. Facts On File, New York, NY, USA, revised edition, 2008. ISBN 0-8160-6695-7 (hardcover). xv + 368 pp. LCCN ???? URL http:

//www.loc.gov/catdir/toc/
ecip0720/2007023966.html.

Zenil:2013:CUU

[Zen13] Hector Zenil, editor. A Computable Universe: Understanding and Exploring Nature as Computation. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4374-29-6. xliv + 810 pp. LCCN QA267.7 .C676 2013. URL http:/

/www.worldscientific.com/

                                        worldscibooks/10.1142/8306.

Foreword by Roger Penrose.